# Computation method for perfectoid purity and perfectoid BCM-regularity

### v5

@@ 6. Case of graded rings
Let $S=\bigoplus_{i\in\mathbb{Q}_{\geq 0}}S_{i}$ be a Noetherian graded ring such that $S_{0}$ is a divisorial valuation ring with $p$ in the maximal ideal of $S_{0}$, $S$ is $p$-torsion free, and $S_{i}=0$ if $i\notin\mathbb{Z}$. We set $\mathfrak{m}:=(p,S_{>0})$, $R:=S_{\mathfrak{m}}$, and $d:=\dim{\overline{R}}$. Then $R$ satisfies conditions in Notation 5.1. Furthermore, we define $S^{(m)}:=\bigoplus_{i\geq 0}S_{im}$ as graded ring for every $m\in\mathbb{Q}_{>0}$. For a graded $S$-module $M$ and $m\in\mathbb{Q}$, we define a graded $S$-module $M(m)$ by $M(m)_{i}=M_{i+m}$ for all $i\in\mathbb{Q}$. Since $H^{d}_{\mathfrak{m}}(\overline{R})\simeq H^{d}_{\mathfrak{m}}(\overline{S})$, we endow $H^{d}_{\mathfrak{m}}(\overline{R})$ with a graded structure. We set $a(\overline{S}):=\max\{m\in\mathbb{Z}\mid H^{d}_{\mathfrak{m}}(\overline{S})_{m}\neq 0\}$. Furthermore, we define the $\mathbb{Q}$-graded structure on $W_{n}(S)$ by
@@ 6.1. Case of $a(\overline{S})<0$
We use the notation introduced in Notation 6.1 and assume $R$ is quasi-$F$-split. Let $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})\backslash\{0\}$ be a homogeneous element of degree $\nu$. Then we can construct sequences $\{\eta_{i}\in H^{d}_{\mathfrak{m}}(\overline{R})\backslash\{0\}\}_{i\geq 0}$ and $\{n_{i}\in\mathbb{Z}_{\geq 1}\}_{i\geq 0}$ using a procedure in Definition 5.4 such that $\eta_{i}$ is a homogeneous element of degree $p^{\sum_{j=0}^{i-1}n_{j}}\nu$.
@@
(Theorem C) We use the notation introduced in Notation 6.1. We assume $\overline{S}$ is normal quasi-Gorenstein and satisfies $a(\overline{S})<0$. We further assume that $\operatorname{Spec}{\overline{S}}\backslash\{\mathfrak{m}\overline{S}\}$ is strongly $F$-regular. If $R$ is quasi-$F$-split, then $R$ is perfectoid BCM-regular. Furthermore, if $p=2$, the converse implication holds.
@@
Let $k$ be an algebraically closed field of characteristic $p$. Let $X$ be a normal integral projective flat scheme over $\operatorname{Spec}{W(k)}$. We assume $-K_{X}$ is ample $\mathbb{Q}$-Cartier. We set
@@
and $-K_{\overline{X}}$ is ample, we obtain $a(\overline{S^{\prime}})<0$. Furthermore, since $\overline{X}$ is strongly $F$-regular, so is $\operatorname{Spec}{S^{\prime}}\backslash\{S^{\prime}_{>0}\}$. Therefore, it is enough to show that $S/pS\simeq S^{\prime}$ as graded rings by Theorem 6.3. We note that if $S/pS\simeq S^{\prime}$, then we have
@@
Let $k$ be an algebraically closed field of characteristic $p$. Let $X$ be a normal integral projective flat scheme over $\operatorname{Spec}{W(k)}$ and $\Delta$ an effective $\mathbb{Q}$-Weil divisor on $X$, whose components are flat over $\operatorname{Spec}{W(k)}$. We assume that $-(K_{X}+\Delta)$ is ample $\mathbb{Q}$-Cartier and $\Delta$ has standard coefficients. We set
@@
Indeed, the cone of $(\overline{X},\overline{\Delta})$ coincides with $\overline{S}$ by (a) and (b), the $a$-invariant $a(\overline{S})$ is negative by (c). By an argument in the proof of Corollary 6.4, we have that $X$ is divisorially Cohen-Macaulay, that is, for all divisorial sheaves on $X$ are Cohen-Macaulay. We set $\Delta:=\sum a_{i}E_{i}$, where each $E_{i}$ is a prime divisor and $a_{i}$ has a form $a_{i}=(m_{i}-1)/m_{i}$ for some integer $m_{i}\geq 2$. Since $(\overline{X},\overline{\Delta})$ is strongly $F$-regular, we have $\lfloor\overline{\Delta}\rfloor=0$, and in particular, we have $E_{i}|_{\overline{X}}$ is a reduced divisor. Therefore, we obtain that $\mathcal{O}_{X}(-m(K_{X}+\Delta))\otimes_{\mathcal{O}_{X}}\mathcal{O}_{\overline{X}}\simeq\mathcal{O}_{\overline{X}}(-m(K_{\overline{X}}+\overline{\Delta}))$ on codimension one points of $\overline{X}$. Since $X$ is divisorially Cohen-Macaulay, we obtain condition (a). Next, we prove the holding condition (b). We take an integer $m\geq 0$ and set $L:=\lfloor-m(K_{X}+\Delta)\rfloor$, which is $\mathbb{Q}$-Cartier since $\overline{X}$ is $\mathbb{Q}$-factorial. Since the Picard rank of $\overline{X}$ is one and $-K_{X}$ is ample, we obtain that $L-K_{X}$ or $-L$ is ample. By [KTTWYY1]*Theorem 5.1, we obtain condition (b) if $\dim{\overline{X}}\geq 2$. Therefore, it is enough to show condition (c). By the Serre duality, we obtain
@@
Since $-(K_{\overline{X}}+\overline{\Delta})$ is ample, it is enough to show that $\lceil m\overline{\Delta}\rceil\leq(m+1)\overline{\Delta}$ for every integer $m\geq 0$. We take a prime divisor $E$ on $\overline{X}$ and a positive integer $n$ with $\frac{n-1}{n}=\mathrm{ord}_{E}(\overline{\Delta})$. We take integers $q\geq 0$ and $n>r\geq 0$ such that $m=nq+r$. Then we have
@@
is perfectoid BCM-regular. On the other hand, we see that $S=W(k)[x,y,z]/(z^{2}+x^{3}+y^{5})$, thus $S/pS$ is not strongly $F$-regular by [Hara2]*Theorem 1.1. Therefore, the ring $S$ satisfies the desired conditions. Next, we assume $p\geq 7$. By [CTW18]*Theorem 1.1, there exists a projective klt surface $\overline{X}$ such that $-K_{\overline{X}}$ is ample and $\overline{X}$ is not globally $F$-regular. By [CTW18]*Theorem 1.1 and [ABL]*Proposition 2.5, a log resolution of $\overline{X}$ lifts to $W(k)$. By [BBKW]*Proposition 6.2, there exists a flat normal integral projective scheme $X$ over $\operatorname{Spec}{W(k)}$ such that $X\times_{\operatorname{Spec}{W(k)}}\operatorname{Spec}{k}\simeq\overline{X}$ and $K_{X}$ is $\mathbb{Q}$-Cartier. Since $-K_{\overline{X}}$ is ample, so is $-K_{X}$. By Corollary 6.4, the localization $R$ of
@@ 6.2. Case of $a(\overline{S})\geq 0$
We use the notation introduced in Notation 6.1. We assume $R$ is Gorenstein and $a(\overline{S})=0$. Then $\operatorname{ht}(R)=\operatorname{ht}(\overline{R})$
@@
By Proposition 4.6, we have $\operatorname{ht}(R)\leq\operatorname{ht}(\overline{R})$. We set $n:=\operatorname{ht}(R)$, then it is enough to show that $\overline{R}$ is $n$-quasi-$F$-split. If $n=1$, the assertion follows from $Q_{R,1}\simeq Q_{\overline{R},1}\simeq F_{*}\overline{R}$, thus we assume $n\geq 2$. Let $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ be a homogeneous non-zero element of degree zero, then $\eta$ is contained in the socle of $H^{d}_{\mathfrak{m}}(\overline{R})$ by $a(\overline{S})=0$. We consider the following commutative diagram;
[…]
Since $\Phi_{R,n}(\eta)\neq 0$, $\Phi_{R,n-1}(\eta)=0$, and $R$ is Cohen-Macaulay, there uniquely exists $\eta^{\prime}\in H^{d}_{\mathfrak{m}}(\overline{R})$ such that $\Phi_{R,n}(\eta)=V^{n-1}(\eta^{\prime})$. By Proposition 6.2, the degree of $\eta^{\prime}$ is zero. By the exact sequence
$0\to F^{n-1}_{*}B_{\overline{R}}\to Q_{\overline{R},n}\to F_{*}\overline{R}\to 0$
and the Cohen-Macaulayness of $\overline{R}$, the homomorphism $V^{n-1}\colon H^{d}_{\mathfrak{m}}(F^{n-1}_{*}B_{\overline{R}})\to H^{d}_{\mathfrak{m}}(Q_{\overline{R},n})$ is injective, thus it is enough to show that $\eta^{\prime}$ is not contained in the kernel of $\alpha$. By the exact sequence
@@
[…]
@@
We take a surjection $S^{\prime}:=\mathbb{Z}_{p}[[x,y,z,x^{\prime},y^{\prime},z^{\prime}]]\to R^{\prime}$ and we define perfectoids $S^{\prime}_{\infty}$ and $R^{\prime}_{\infty}:=R^{S^{\prime}_{\infty}}_{perfd}$ as in [p-pure]*Section 4.3. By the natural homomorphism $H^{3}_{\mathfrak{m}}(R)\otimes_{\mathbb{Z}_{p}}H^{3}_{\mathfrak{m}}(R)\to H^{5}_{\mathfrak{n}}(R^{\prime})$, the image of the element $[z^{2}/xyp]\otimes[z^{\prime 2}/x^{\prime}y^{\prime}p]$ is
$[z^{2}z^{\prime 2}/xyx^{\prime}y^{\prime}p],$
@@ Appendix A Construction of functorial test perfectoids
First, we construct a perfectoid $T(R)$ over a $\mathbb{Z}_{(p)}$-algebra $R$. We set $A:=\mathbb{Z}_{p}[X]^{\wedge(p,d)}$ and $d:=X-p$, and we define the $\mathbb{Z}_{p}$-algebra homomorphism $\phi\colon A\to A$ by $\phi(X)=X^{p}$, then $(A,d)$ is a prism. We define $A_{\infty}$ by the $(p,d)$-adic completion of perfection of $A$. Let $A\{R\}:=A\{X_{f}\mid f\in R\}$ be a free $\delta$-ring as in [BS]*Lemma 2.11, then $(A\{R\}^{\wedge(p,d)},d)$ is a prim. Taking a perfection and $(p,d)$-completion, we obtain a perfect prism $(A\{R\}_{\infty},(d))$. We set $Z\{R\}_{\infty}:=A\{R\}_{\infty}/(d)$, then we have the $\mathbb{Z}_{(p)}$-algebra homomorphism $\mathbb{Z}_{(p)}[R]:=\mathbb{Z}_{(p)}[X_{f}\mid f\in R]\to A\{R\}_{\infty}$ defined by $X_{f}\mapsto X_{f}$. We set

### v7

@@ 6. Case of graded rings
Let $S=\bigoplus_{i\in\mathbb{Q}_{\geq 0}}S_{i}$ be a Noetherian graded ring such that $S_{0}$ is a divisorial valuation ring with $p$ in the maximal ideal of $_{0}, SS$ is $p$-torsion free, and $S_{i}=0$ if $i\notin\mathbb{Z}$. We set $\mathfrak{m}:=(p,S_{>0})$, $R:=S_{\mathfrak{m}}$, and $d:=\dim{\overline{R}}$. Then $R$ satisfies conditions in Notation 5.1. Furthermore, we define $S^{(m)}:=\bigoplus_{i\geq 0}S_{im}$ as graded ring for every $m\in\mathbb{Q}_{>0}$. For a graded $S$-module $M$ and $m\in\mathbb{Q}$, we define a graded $S$-module $M(m)$ by $M(m)_{i}=M_{i+m}$ for all $i\in\mathbb{Q}$. Since $H^{d}_{\mathfrak{m}}(\overline{R})\simeq H^{d}_{\mathfrak{m}}(\overline{S})$, we endow $H^{d}_{\mathfrak{m}}(\overline{R})$ with a graded structure. We set $a(\overline{S}):=\max\{m\in\mathbb{Z}\mid H^{d}_{\mathfrak{m}}(\overline{S})_{m}\neq 0\}$. Furthermore, we define the $\mathbb{Q}$-graded structure on $W_{n}(S)$ by
@@ 6.1. Case of $a(\overline{S})<0$
We use the notation introduced in Notation 6.1 and assume $R$ is quasi-$F$-split. Let $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})\backslash\{0\}$ be a homogeneous element of degree $\nu$. Then we can construct sequences $\eta_{i}\in H^{d}_{\mathfrak{m}}(\overline{R})\backslash\{0\}\}_{i\geq 0}$ and $\{n_{i}\in\mathbb{Z}_{\geq 1}\}_{i\geq 0}$ using a procedure in Definition 5.4 such that $\eta_{i}$ is a homogeneous element of degree $p^{\sum_{j=0}^{i-1}n_{j}}\nu$.
@@
[…]
@@
Let $k$ be an algebraically closed field of characteristic $p$. Let $X$ be a normal integral projective flat scheme over $\operatorname{Spec}{W(k)}$. We assume $K_{X}$ is ample $\mathbb{Q}$-Cartier. We set
@@
[…]
@@
Let $k$ be an algebraically closed field of characteristic $p$. Let $X$ be a normal integral projective flat scheme over $\operatorname{Spec}{W(k)}$ and $\Delta$ an effective $\mathbb{Q}$-Weil divisor on $X$, whose components are flat over $\operatorname{Spec}{W(k)}$. We assume that $(K_{X}+\Delta)$ is ample $\mathbb{Q}$-Cartier and $\Delta$ has standard coefficients. We set
@@
Indeed, the cone of $(\overline{X},\overline{\Delta})$ coincides with $\overline{S}$ by (a) and (b), the $a$-invariant $a(\overline{S})$ is negative by (c). By an argument in the proof of Corollary 6.4, we have that $X$ is divisorially Cohen-Macaulay, that is, for all divisorial sheaves on $X$ are Cohen-Macaulay. We set $\Delta:=\sum a_{i}E_{i}$, where each $E_{i}$ is a prime divisor and $a_{i}$ has a form $a_{i}=(m_{i}-1)/m_{i}$ for some integer $m_{i}\geq 2$. Since $(\overline{X},\overline{\Delta})$ is strongly $F$-regular, we have $\lfloor\overline{\Delta}\rfloor=0$, and in particular, we have $E_{i}|_{\overline{X}}$ is a reduced divisor. Therefore, we obtain that $\mathcal{O}_{X}(-m(K_{X}+\Delta))\otimes_{\mathcal{O}_{X}}\mathcal{O}_{\overline{X}}\simeq\mathcal{O}_{\overline{X}}(-m(K_{\overline{X}}+\overline{\Delta}))$ on codimension one points of $\overline{X}$. Since $X$ is divisorially Cohen-Macaulay, we obtain condition (a). Next, we prove the holding condition (b). We take an integer $m\geq 0$ and set $L:=\lfloor-m(K_{X}+\Delta)\rfloor$, which is $\mathbb{Q}$-Cartier since $\overline{X}$ is $\mathbb{Q}$-factorial. Since the Picard rank of $\overline{X}$ is one and $K_{X}$ is ample, we obtain that $L-K_{X}$ or L-L is ample. By [KTTWYY1]*Theorem 5.1, we obtain condition (b) if $\dim{\overline{X}}\geq 2$. Therefore, it is enough to show condition (c). By the Serre duality, we obtain
@@
Since $(K_{\overline{X}}+\overline{\Delta})$ is ample, it is enough to show that $\lceil m\overline{\Delta}\rceil\leq(m+1)\overline{\Delta}$ for every integer $m\geq 0$. We take a prime divisor $E$ on $\overline{X}$ and a positive integer $n$ with $\frac{n-1}{n}=\mathrm{ord}_{E}(\overline{\Delta})$. We take integers $q\geq 0$ and $n>r\geq 0$ such that $m=nq+r$. Then we have
@@
is perfectoid BCM-regular. On the other hand, we see that $S=W(k)[x,y,z]/(z^{2}+x^{3}+y^{5})$, thus $S/pS$ is not strongly $F$-regular by [Hara2]*Theorem 1.1. Therefore, the ring $S$ satisfies the desired conditions. Next, we assume $p\geq 7$. By [CTW18]*Theorem 1.1, there exists a projective klt surface $\overline{X}$ such that $K_{\overline{X}}$ is ample and $\overline{X}$ is not globally $F$-regular. By [CTW18]*Theorem 1.1 and [ABL]*Proposition 2.5, a log resolution of $\overline{X}$ lifts to $W(k)$. By [BBKW]*Proposition 6.2, there exists a flat normal integral projective scheme $X$ over $\operatorname{Spec}{W(k)}$ such that $X\times_{\operatorname{Spec}{W(k)}}\operatorname{Spec}{k}\simeq\overline{X}$ and $K_{X}$ is $\mathbb{Q}$-Cartier. Since $K_{\overline{X}}$ is ample, so is $K_{X}$. By Corollary 6.4, the localization $R$ of
@@ 6.2. Case of $a(\overline{S})\geq 0$
We use the notation introduced in Notation 6.1. Assume that $R$ is Gorenstein and that $a(\overline{S})=0$. Then $\operatorname{ht}(R)=\operatorname{ht}(\overline{R})$.
@@
By Proposition 4.6, we have $\operatorname{ht}(R)\leq\operatorname{ht}(\overline{R})$. Let $n:=\operatorname{ht}(R)$. Then it suffices to show that $\overline{R}$ is $n$-quasi-$F$-split. If $n=1$, the assertion follows immediately from the isomorphisms $Q_{R,1}\simeq Q_{\overline{R},1}\simeq F_{*}\overline{R}$. Hence, we may assume $n\geq 2$.
Let $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ be a homogeneous nonzero element of degree zero. Since $a(\overline{S})=0$, the socle of $H^{d}_{\mathfrak{m}}(\overline{R})$ is concentrated in degree zero, and thus $\eta$ lies in the socle.
We consider the following commutative diagram:
${H^{d}_{\mathfrak{m}}(\overline{R})}$${H^{d}_{\mathfrak{m}}(Q_{R,n})}$${H^{d}_{\mathfrak{m}}(F^{n}_{*}\overline{R})}$${H^{d}_{\mathfrak{m}}(Q_{\overline{R},n})}$${H^{d}_{\mathfrak{m}}(F^{n-1}_{*}B_{\overline{R}})}$$\scriptstyle{\Phi_{R,n}}$$\scriptstyle{\Phi_{\overline{R},n}}$$\scriptstyle{V^{n-1}}$$\scriptstyle{\alpha}$$\scriptstyle{V^{n-1}}$
Since $\Phi_{R,n}(\eta)\neq 0$ and $\Phi_{R,n-1}(\eta)=0$, and $R$ is Cohen–Macaulay, there exists a unique $\eta^{\prime}\in H^{d}_{\mathfrak{m}}(\overline{R})$ such that $\Phi_{R,n}(\eta)=V^{n-1}(\eta^{\prime})$. By Proposition 6.2, the degree of $\eta^{\prime}$ is zero.
Moreover, since $\overline{R}$ is not $F$-pure, the map $H^{d}_{\mathfrak{m}}(\overline{R})_{0}\to H^{d}_{\mathfrak{m}}(F_{*}\overline{R})_{0}$ is the zero map. In particular, $\eta^{\prime}$ is not contained in the image of Frobenius. Thus, under the exact sequence
@@
the element $\eta^{\prime}$ does not map to zero under the induced map $\alpha$ on local cohomology. Hence, to prove that $\eta^{\prime}$ survives in $H^{d}_{\mathfrak{m}}(Q_{\overline{R},n})$, it suffices to show that the map
$V^{n-1}_{0}\colon H^{d}_{\mathfrak{m}}(F^{n-1}_{*}B_{\overline{R}})_{0}\to H^{d}_{\mathfrak{m}}(Q_{\overline{R},n})_{0}$
is injective.
We now prove that $F_{0}\colon H^{d}_{\mathfrak{m}}(W_{m}(\overline{R}))_{0}\to F_{*}H^{d}_{\mathfrak{m}}(W_{m}(\overline{R}))_{0}$ is the zero map for every $1\leq m\leq n-1$ by induction on $m$.
For the base case $m=1$, the assertion follows from the fact that $\overline{R}$ is not $F$-pure. Assume the claim holds for m1m-1 with $m\leq n-1$.
Since $\overline{R}$ is not (n1)(n-1)-quasi-$F$-split, the composition
$H^{d}_{\mathfrak{m}}(W_{m}(\overline{R}))_{0}\xrightarrow{F}F_{*}H^{d}_{\mathfrak{m}}(W_{m}(\overline{R}))_{0}\to H^{d}_{\mathfrak{m}}(Q_{\overline{R},m})_{0}$
is zero.
From the short exact sequence
$0\to F_{*}W_{m-1}(\overline{R})\xrightarrow{VF}F_{*}W_{m}(\overline{R})\to Q\to 0,$
and using the induction hypothesis that $F_{0}$ on $H^{d}_{\mathfrak{m}}(W_{m-1}(\overline{R}))_{0}$ is zero, we find that the map $F_{*}H^{d}_{\mathfrak{m}}(W_{m}(\overline{R}))_{0}\to H^{d}_{\mathfrak{m}}(Q_{\overline{R},m})_{0}$ is injective. Thus, $F_{0}$ on $H^{d}_{\mathfrak{m}}(W_{m}(\overline{R}))_{0}$ must be zero.
Next, consider the following commutative diagram, where each row is a short exact sequence:
${0}$${H^{d}_{\mathfrak{m}}(F^{n-2}_{*}\overline{R})_{0}}$${H^{d}_{\mathfrak{m}}(W_{n-1}(\overline{R}))_{0}}$${H^{d}_{\mathfrak{m}}(W_{n-2}(\overline{R}))_{0}}$${0}$${H^{d}_{\mathfrak{m}}(F^{n-2}_{*}B_{\overline{R}})_{0}}$${H^{d}_{\mathfrak{m}}(B_{\overline{R},n-1})_{0}}$${H^{d}_{\mathfrak{m}}(B_{\overline{R},n-2})_{0}}$${0.}$$\scriptstyle{V^{n-2}}\scriptstyle{\cong}\scriptstyle{\cong}\scriptstyle{\cong}$$\scriptstyle{V^{n-2}}$
By the above, each vertical map is an isomorphism. Therefore, we obtain the injection:
$H^{d}_{\mathfrak{m}}(F^{n-2}_{*}B_{\overline{R}})_{0}\hookrightarrow H^{d}_{\mathfrak{m}}(B_{\overline{R},n-1})_{0}.$
Furthermore, from the exact sequence
$0\to B_{\overline{R},n-1}\xrightarrow{V}Q_{\overline{R},n}\to\overline{R}\to 0,$
and the fact that $\overline{R}$ is Cohen–Macaulay, we deduce that the map
$H^{d}_{\mathfrak{m}}(B_{\overline{R},n-1})\to H^{d}_{\mathfrak{m}}(Q_{\overline{R},n})$
is injective.
Combining the above, we obtain the desired injectivity:
$V^{n-1}_{0}\colon H^{d}_{\mathfrak{m}}(F^{n-1}_{*}B_{\overline{R}})_{0}\hookrightarrow H^{d}_{\mathfrak{m}}(Q_{\overline{R},n})_{0}.$
This completes the proof. ∎
@@
We take a surjection $S^{\prime}:=\mathbb{Z}_{p}[[x,y,z,x^{\prime},y^{\prime},z^{\prime}]]\to R^{\prime}$ and we define perfectoids $S^{\prime}_{\infty}$ and $R^{\prime}_{\infty}:=R^{S^{\prime}_{\infty}}_{perfd}$ as in [p-pure]*Section 4.3. By the natural homomorphism $H^{3}_{\mathfrak{m}}(R)\otimes_{\mathbb{Z}_{p}}H^{3}_{\mathfrak{m}}(R)\to H^{5}_{\mathfrak{n}}(R^{\prime})$, the image of the element $z^{2}/xyp]\otimes[z^{\prime 2}/x^{\prime}y^{\prime}p]$ is
$z^{2}z^{\prime 2}/xyx^{\prime}y^{\prime}p],$
@@ Appendix A Construction of functorial test perfectoids
First, we construct a perfectoid $T(R)$ over a $\mathbb{Z}_{(p)}$-algebra $R$. We set $A:=\mathbb{Z}_{p}[X]^{\wedge(p,d)}$ and d:=Xpd:=X-p, and we define the $\mathbb{Z}_{p}$-algebra homomorphism $\phi\colon A\to A$ by $\phi(X)=X^{p}$, then $(A,d)$ is a prism. We define $A_{\infty}$ by the $(p,d)$-adic completion of perfection of $A$. Let $A\{R\}:=A\{X_{f}\mid f\in R\}$ be a free $\delta$-ring as in [BS]*Lemma 2.11, then $(A\{R\}^{\wedge(p,d)},d)$ is a prim. Taking a perfection and $(p,d)$-completion, we obtain a perfect prism $(A\{R\}_{\infty},(d))$. We set $Z\{R\}_{\infty}:=A\{R\}_{\infty}/(d)$, then we have the $\mathbb{Z}_{(p)}$-algebra homomorphism $\mathbb{Z}_{(p)}[R]:=\mathbb{Z}_{(p)}[X_{f}\mid f\in R]\to A\{R\}_{\infty}$ defined by $X_{f}\mapsto X_{f}$. We set